OpenStudy (anonymous):

The length of the top of a table is 4m greater than the width. The area is 96m^2. Find the dimensions of the table. Width= Length=

OpenStudy (anonymous):

so let the width = x, therefore length = x+4

OpenStudy (anonymous):

Area is length times width, so area = x*(x+4) = 96

OpenStudy (anonymous):

Now solve for x and you've got your answer

OpenStudy (anonymous):

Is width = 92m^2 Length= 96m^2

OpenStudy (anonymous):

no, try solving \[x^2+4x-96=0\]

OpenStudy (anonymous):

(x-8)(x+12)

OpenStudy (anonymous):

correct, now which value of x is the logical answer?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

your equation should read (x-8)(x+12)=0 the value of 12 does not make this equation balanced

OpenStudy (anonymous):

x=-12 or x=8

OpenStudy (anonymous):

so which root is logical? is it possible to have -12 width?

OpenStudy (anonymous):

No x=8

OpenStudy (anonymous):

so width is 8, length is 8+4 = 12 Always check, 8 times 12 = 96

OpenStudy (anonymous):

Thanx